ANALYSIS OF PHOSPHINE IN THE 3 μm REGION FOR PLANETARY APPLICATIONS.

\underline{R. A. H. BUTLER}, L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109,USA; I. KLEINER and L. SAGUI, Laboratoire de Photophysique Moléculaire, Université Paris Sud, 91405 Orsay Cedex, France.

Progress on the analysis of the high resolution spectrum of PH_{3} in the 3 μm region between 3100 to 3600 cm^{−1} is reported. The objective is to provide line parameters for remote sensing of Saturn. For this, laboratory spectra have been recorded at a resolution of 0.012 cm^{−1} using the Fourier transform spectrometer located at the Kitt Peak Observatory. Using ground state combination differences, several hundred transitions have been assigned in ν_{1}+ν_{2}, ν_{3}+ν_{4}, ν_{1}+ν_{4} and ν_{2}+2ν_{4} centered around 3307, 3425, 3432 and 3215 cm^{−1} respectively. Analysis and preliminary modeling of the interacting energy states will be presented. ^{a}

aPart of the research reported in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.